Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB SQS SXP PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF not2MigrateOrArchive MigratedByGroup
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 384     Entry time: 09 Jan 2023, 10:10
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 09.01.2023 
General Announcement:
- Interlock tests light taking place today and tomorrow, starting at Desy premises and tunnels.
- On 23.01 there is the Interlock tests at XHEXP1. 
- Zunaira A. asks for one experimental hutch in all SASEs (optics hutch can also be) to be closed for tomorrow to check some signals for tunnels. Hutches need to be closed at around 09:30-10:00. 
SQS accepts to interlock the hutch. MID also says it is possible but a precise time needs to be communicated.
- On 17-18.January the 5Hz operation will happen. Needed groups are already informed and tests shall be transparent for the instruments. Tim Wilksen will prepare a schedule to be distributed as soon as 
available.

CTRL:
- Influx DB: two nodes are available. A 6-month policy notice is the one to be used by Karabo. An e-mail will be sent.

DA:
- Nothing to report.
EEE:
- SA2 lost power on the 31.01. The CPUs in the instruments were not affected; the CPU in SA2 laser was affected and parameters had to be reset. Impression is that UPS did not act. EEE needs more 
information.
- A. Galler says that it relates to the broken UPS in SA2 area. A possible automatic check run/done by MPS triggered the emergency OFF. A new battery for the UPS is ordered and it is to be delivered in 
future.

VAC:
- All relevant tasks for WMP are closed.

XPD:
- XPD would like to continue in XTD9 and XTD6 after the interlock tests to conclude minor activities.

DET:
- Nothing to report.

ITDM: 
- Reminder to all instruments and support groups: finish accessing and closing the proposals (deadline has already passed) before they will be discarded.

XRO:
- Will need access to tunnels XTD6 at late January. The access is not related to test components.

LAS:
- Bringing lasers back up. In SA2 there was a rough start due to the power OFF but seems all in good track.

UNSYS:
- SA3 undulators: rotary encoders could be used to control the undulator gaps.
- Suren K. informs that partially recovered encoders for the 3 upstream segments. Spare parts for the next two devices will arrive in some months. Root cause for the SA3 power issue is not understood. 
Undulators can be operated.
- Power distribution terminals in SA2 (2 units) and 1 unit in SA1 needed replacements.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- Request to support groups: if external companies need access to hutch, please remind these companies to do not use dirty shoes when accessing any hutch.

HED:
- Nothing to report.

SCS:
- Issue with XGM gas supply last week (valve was closed by someone unknown). Check functionality is ongoing but all seems to be fine. 
- Transport taking place may affect access in areas around SCS, including XTD10.

SQS:
- Nothing to report.
ELOG V3.1.4-7c3fd00